# Topolectrical circuits – recent experimental advances and developments

### v1

@@ IV.2 Floquet circuits and the observation of temporal phenomena
Typical realizations of tight-binding condensed matter phenomena map the lattice model onto a real space physical array, and is often met with challenges in implementing or maintaining a large number of lattice sites to within acceptance levels of error tolerance. One promising alternative approach is to instead realize the model in the frequency lattice, through a time-periodic driving approach known as Floquet engineering [367, 368, 369, 370, 371, 372, 373, 374, 375, 376, 377, 378, 379, 380, 381, 382, 383, 384]. In analogy to the Fourier correspondence between a periodic Brillouin zone and its real-space lattice, a system that is periodically driven in time should also correspond to a lattice in frequency space. This powerful analogy enables the implementation of complicated hoppings and unit cells in terms of appropriately programmed temporal driving frequency components, all within a single (time-periodic) physical unit cell.
@@
A key difference between Floquet-driven Hamiltonian systems and Floquet-driven electrical circuits is that in the latter (electrical circuits), the main physics is contained in the steady-state relationship between the current and the voltage i.e. the circuit Laplacian that encapsulates Kirchhoff’s law (refer to Section II.1). As such, impedance measurements probe the properties of the Laplacian , rather than the time-evolution generator . As explained in Ref. [101], this fundamental difference is reflected in the frequency dependence of the Floquet operator: while ordinary Floquet Hamiltonians can only contain a linear “background” potential in the frequency lattice [385, 386], a Floquet Laplacian can depend on the frequency in arbitrarily complicated manners, depending on the frequency dependencies of its constituent components i.e. or for capacitive and inductive components. As such, Floquet circuit Laplacians can be engineered to possess a variety of background “potentials” in the frequency lattice, such that one can also probe impurity or edge phenomena i.e. topological edge states in appropriately designed temporally driven electrical circuits.
@@
Stegmaier et al. [101] designed an a circuit with two logical nodes that assumes the topological SSH form in the frequency domain [Fig. 7a1]. By employing a combination of AD633 analog multipliers and capacitors [Fig. 7a2], effectively time-varying capacitors were achieved, modulated at far higher frequencies (order of MHz) than what is possible with mechanical means. To recreate an SSH model in frequency space, the inter- and intra-node capacitances are time-modulated such that their temporal functional dependence resembles that of the momentum dependence of the conventional SSH model. Notably, the Floquet driving protocol is such that the effective background potential diverges at the zeroth “site” in the frequency lattice, such that it acts as a strong “barrier”. As conspicuously measured, topology boundary states exists at frequencies very near these barriers when the effective SSH parameters are tuned to topologically nontrivial values [Fig. 7a3]. These barrier states exhibit localized profiles in frequency space, as can be obtained by Fourier transforming the measured temporal voltage oscillations [Fig. 7a4]. Other than showcasing the utility of analog multipliers in effectively time-modulating circuit components at very high frequencies, this work [Ref. [101]] also demonstrated that Floquet circuit Laplacians can exhibit highly customizable inhomogeneities in the frequency lattice, an important realization that has far-reaching implications in topological signal processing.
@@
Time-modulated electrical circuits can also be used to demonstrate dynamical phenomena such as Thouless pumping [388]. Traditionally, the Thouless pump describes the adiabatic, quantized transport of charge across a bulk gap [389], as can be deduced from spectral flow arguments. In the regime of classical electrical circuits, charge quantization is not accessible, but spectral flow can still be measured since the Laplacian spectrum corresponds to impedance resonances. Stegmaier et al. [102] demonstrated such spectral pumping in an electrical circuit with inductors that are effectively time-modulated with analog multipliers [Fig. 7c1]. As in Ref. [101], the time modulation is controlled by an external voltage signal which can be customized at will – here, the modulation is chosen to yield the Aubry-André-Harper (AAH) model. By measuring the impedances within the circuit, eigenfrequency bands were reconstructed as a function of the periodically pumped phase. Indeed, the expected AAH spectral flow was observed under PBCs [Fig. 7c2], and topological in-gap states can also be seen under OBCs [Fig. 7c3]. It is important to note that, unlike the quantized charge transport predicted by the Kubo formula in quantum settings, here there is no occupied ground state, and the observed state pumping is purely that of the spectral band resonances.
@@
Another noteworthy result in Ref. [387] is the demonstration of the generalized Brillouin zone using the Laplace transform rather than the Fourier transform applied to the measured voltage distribution. The non-Hermiticity renders the wavevector complex, making it infeasible to obtain the Brillouin zone directly from the Fourier transform. However, Wu et al. utilized the Laplace transform method to effectively recover the generalized Brillouin zone, where different onsite admittances correspond to different contours of the band structure. Such site-resolved field distributions are highly accessible in topolectrical circuits, which is crucial for studying higher-dimensional lattices. For example, in another work [391], Wu et al. observed degeneracy conversions, including Weyl point–nodal line conversion and nodal line–entangled nodal line conversion.
@@ IV.3 Hyperbolic lattices and their realizations in electrical circuits
Lately, there has been much interesting in the band structure and topological properties of hyperbolic lattices, which can be regarded as a vast generalization of regular lattices in negatively curved space [395, 396, 397, 393, 398, 399, 400, 401, 402, 403, 404, 405]. In 2D Euclidean (flat) space, regular -gons can be tessellated such that every vertex contains branches, if . This leaves only three possibilities corresponding to the triangular, square and hexagonal lattices respectively. Relaxing this constraint to results in an infinite variety of so-called hyperbolic lattices, which however need to be embedded in negatively curved space 111The opposite constraint gives the five possibilities , respectively corresponding to the five platonic solids i.e. icosahedron, dodecahedron, octahedron, cube and tetrahedron, all tiling the sphere.. While this requirement of negatively curved space embedding makes them challenging to realize in most metamaterial platforms, electrical circuit connections can describe any desired graph structure, and thus emerge as promising physical platforms for hyperbolic lattices. This negative curvature fundamentally alters the propagation of modes, enriching the behavior of fundamental phenomena such as topological propagation [407]. Due to the relative proliferation of boundary terminations, a hyperbolic lattice structure also alters the nature of the non-Hermitian bulk-boundary correspondences in the presence of asymmetric lattice hoppings [408]. The ability of electrical circuits to simulate lattices with high coordination numbers is also helpful in the realization of higher dimensional lattices [409, 410, 411, 114, 50, 129, 105, 130].
@@
where the phase is controlled by adjusting the external voltages of the analog multipliers. The diagonal entries introduce a constant offset in the admittance spectrum, while the off-diagonal terms generate the controllable complex phase between ports and .
A key feature of a finite hyperbolic lattices is that the number boundary sites scales exponentially with the radius (i.e. number of lattice generations), reminiscent of fractal lattices [412]. As such, unlike in Euclidean lattices, boundary sites can dominate the Hilbert space and greatly alter bulk-boundary correspondences, even in Hermitian setups. Zhang et al. [393] demonstrated the exponential proliferation of zero modes in a hyperbolic circuit [Fig. 9c]. These topological zero modes give rise to a large number of impedance resonances at the edge nodes, while no resonances occur in the bulk nodes.
This boundary sensitivity is reflective of the small-world connectivity of a hyperbolic lattice, a property that have inspired their application in holographic duality [413, 414, 415, 416, 417, 418, 419]. The high connectivity also gives rise to heightened susceptibility to lattice defects or perturbations. Pei et al. [394] experimentally constructed a hyperbolic Haldane circuit and showed that a single defect in the bulk can induce boundary-dominated one-way propagation. By introducing polygonal bulk defects into the Haldane hyperbolic lattice, robust boundary propagation was observed, even when induced by a single point defect in the bulk [Fig. 9d]. While the presence of zero modes was verified through impedance responses, the one-way propagation of topological edge states was demonstrated using dynamical voltage analysis.
@@ IV.4 Simulations of other topological and quantum systems
Electrical circuits are also very suited for simulating quantum media, both of conventional topological materials and more unconventional setups such as [391, 420, 421, 422]. Described below are some experimental demonstrations, particularly those of models that are hard to simulate on other platforms.
One class of models that are particularly amenable to electrical circuit realizations are the square-roots models. Given a parent Hamiltonian, the square-root of it is a lattice model in which two successive hoppings yield the same connectivity structure as the parent model (up to an unimportant constant shift) [435, 436, 437, 438, 439, 440, 441, 442, 443, 444, 445, 446, 447, 448]. Mathematically, they may look contrived and complicated, but they inherit various porperties from their parent Hamiltonians, such as midgap topological modes [449]. The number of modes, including topological modes, is doubled in these lattices [450]. Electrical circuits are suitable physical platforms due to their versatile connectivity, and Song et al. [423] constructed a 3D square-root higher-order Weyl semi-metal by stacking 2D square-root lattices along the third dimension [Fig. 9a]. From the admittance spectrum, non-zero mid-gap topological hinge states were observed.
@@
Beyond band structures, 2D or 3D electrical circuit arrays can also simulate few-body physics and other phenomena. By mapping the 2-photon degrees of freedom onto a 2D lattice [48, 313, 457], Zhuo et al. [424] simulated interaction-induced flat-band localizations and topological edge states, tuning the effective interaction strength through circuit grounding. By measuring impedance responses and voltage dynamics in the time domain, they were able to resolve two-boson flat bands and topological edge states [Fig. 9c]. On a different note, Yang et al. [425] simulated Wilson fermions – a lattice discretization conventionally used in the context of lattice QCD to avoid fermion doubling – in a circuit lattice, tuning circuit parameters to achieve different Wilson fermion states. Through impedance measurements, they reconstructed the chiral edge domain wall separating two circuits with contrasting fractional Chern numbers, and observed localized corner states arising from two Chern lattices with opposite chiralities [Fig. 9d]. Wang et al. [426] realized the elusive Hopf insulator in a 3D electrical circuit systematically implemented with so-called pseudospin modules. Although this lattice required long-range spin-orbit couplings, the pseudospin modules aided the realization of Hopf bands, including both surface and bulk bands [Fig. 9e].
Electrical circuits have also been used to simulate non-abelian systems, in which physical outcomes depend on the sequence of operations i.e. twists [458, 459, 460, 461]. Various recent experiments have attempted to classically simulate operations that are mathematically equivalent to that experienced by anyonic quasiparticles [462, 463, 464, 465, 466, 467]. Wang et al. [427] demonstrated non-Abelian Anderson localization using a 2D electrical circuit that represents the quasiperiodic AAH model with non-Abelian gauge fields. Both localization and delocalization behaviors were observed through site-resolved impedance spectra and voltage dynamics [Fig. 9f] (see also the experimental realization of non-Abelian inverse Anderson transition in Ref. [468]).
Circuits are also particularly suitable for demonstrating the effects of lattice defects, since node are connected at will. Xie et al. [428] experimentally demonstrated disclination states and their correspondence to real-space topology. Localized topological states at the disclinations—defects in rotational symmetry [469]—were observed in impedance maps and voltage distributions at the lattice nodes [Fig. 9g] (see also Refs. [470, 471]).
@@
Since electrical circuits commonly involve resistive and active elements, they can readily demonstrate how non-Hermiticity interplays with topology without much additional complication. Xie et al. [430] demonstrated the non-Hermitian version of the graphene lattice, which exhibits two non-Hermitian Dirac cones, one experiencing amplification and the other experiencing decay. Site-resolved band measurements reveal unidirectional kink states [480, 481] with dissimilar profiles at the two valleys [Fig. 9j]. Involving asymmetrical real next-nearest-neighbor couplings effectively result in a non-Hermitian single Dirac cone, leading to valley polarization [482, 483, 484, 485] with a large bandwidth.
The introduction of non-Hermiticity can also greatly alter the locality of topological modes. While topological modes are typically strongly confined to the boundaries [30], non-Hermitian skin pumping in the opposite direction can delocalize them while preserving their eigenspectrum, in this case mid-gap zero-eigenvalues [486]. These delocalized topological states, still protected by lattice symmetries, can extend through the bulk, enabling their robustness to be harnessed not only at the boundaries but also within the bulk [487, 488]. Lin et al. [431] demonstrated these extended states in 1D, 2D, and 3D circuits by tuning the non-reciprocal coupling strength [Fig. 9k]. At a critical strength determined by both the topological and skin localization length, the topological modes become extended, a phenomenon which was also studied in Refs. [489, 490, 38, 491].
@@
Topolectrical circuits can simulate not only classical systems, but also quantum setups [496, 497, 498]. By mapping the Hilbert space of few-qubit systems to the nodes of suitably designed circuits, Zou et al. [433] demonstrated [Fig. 9m] the mathematically equivalences of Majorana-like edge states, T-junctions, two-qubit unitary operations, and Grover’s search algorithm. Through their resistor-capacitor-based circuit with segmented fixed resistances, braiding processes can be simulated without significant loss. Zhang et al. [434] demonstrated another example of quantum algorithm simulation using classical circuits. They simulated a two-player quantum game based on AND-OR tree structures [Fig. 9n]. Relays were employed to disconnect each node simultaneously after applying the initial voltage, allowing the system to evolve from the initial state.

### v2

@@ IV.2 Floquet circuits and the observation of temporal phenomena
Typical realizations of tight-binding condensed matter phenomena map the lattice model onto a real space physical array, and are often met with challenges in implementing or maintaining a large number of lattice sites to within acceptance levels of error tolerance. One promising alternative approach is to instead realize the model in the frequency lattice, through a time-periodic driving approach known as Floquet engineering [377, 378, 379, 380, 381, 382, 383, 384, 385, 386, 387, 388, 389, 390, 391, 392, 393, 394]. In analogy to the Fourier correspondence between a periodic Brillouin zone and its real-space lattice, a system that is periodically driven in time should also correspond to a lattice in frequency space. This powerful analogy enables the implementation of complicated hoppings and unit cells in terms of appropriately programmed temporal driving frequency components, all within a single (time-periodic) physical unit cell.
@@
A key difference between Floquet-driven Hamiltonian systems and Floquet-driven electrical circuits is that in the latter (electrical circuits), the main physics is contained in the steady-state relationship between the current and the voltage, i.e., the circuit Laplacian that encapsulates Kirchhoff’s law (refer to Section II.1). As such, impedance measurements probe the properties of the Laplacian , rather than the time-evolution generator . As explained in Ref. [101], this fundamental difference is reflected in the frequency dependence of the Floquet operator: while ordinary Floquet Hamiltonians can only contain a linear “background” potential in the frequency lattice [395, 396], a Floquet Laplacian can depend on the frequency in arbitrarily complicated manner, depending on the frequency dependencies of its constituent components, i.e., or for capacitive and inductive components. As such, Floquet circuit Laplacians can be engineered to possess a variety of background “potentials” in the frequency lattice, such that one can also probe impurity or edge phenomena, i.e., topological edge states in appropriately designed temporally driven electrical circuits.
@@
Stegmaier et al. [101] designed a circuit with two logical nodes that assumes the topological SSH form in the frequency domain [Fig. 7a1]. By employing a combination of AD633 analog multipliers and capacitors [Fig. 7a2], effectively time-varying capacitors were achieved, modulated at far higher frequencies (order of MHz) than what is possible with mechanical means. To recreate an SSH model in frequency space, the inter- and intra-node capacitances are time-modulated such that their temporal functional dependence resembles that of the momentum dependence of the conventional SSH model. Notably, the Floquet driving protocol is such that the effective background potential diverges at the zeroth “site” in the frequency lattice, such that it acts as a strong “barrier”. As conspicuously measured, topology boundary states exist at frequencies very near these barriers when the effective SSH parameters are tuned to topologically nontrivial values [Fig. 7a3]. These barrier states exhibit localized profiles in frequency space, as can be obtained by Fourier transforming the measured temporal voltage oscillations [Fig. 7a4]. Other than showcasing the utility of analog multipliers in effectively time-modulating circuit components at very high frequencies, this work (Ref. [101]) also demonstrated that Floquet circuit Laplacians can exhibit highly customizable inhomogeneities in the frequency lattice, an important realization that has far-reaching implications in topological signal processing.
@@
Time-modulated electrical circuits can also be used to demonstrate dynamical phenomena such as Thouless pumping [398]. Traditionally, the Thouless pump describes the adiabatic, quantized transport of charge across a bulk gap [399], as can be deduced from spectral flow arguments. In the regime of classical electrical circuits, charge quantization is not accessible, but spectral flow can still be measured since the Laplacian spectrum corresponds to impedance resonances. Stegmaier et al. [102] demonstrated such spectral pumping in an electrical circuit with inductors that are effectively time-modulated with analog multipliers [Fig. 7c1]. As in Ref. [101], the time modulation is controlled by an external voltage signal which can be customized at will—here, the modulation is chosen to yield the Aubry-André-Harper (AAH) model. By measuring the impedances within the circuit, eigenfrequency bands were reconstructed as a function of the periodically pumped phase. Indeed, the expected AAH spectral flow was observed under PBCs [Fig. 7c2], and topological in-gap states can also be seen under OBCs [Fig. 7c3]. It is important to note that, unlike the quantized charge transport predicted by the Kubo formula in quantum settings, here there is no occupied ground state, and the observed state pumping is purely that of the spectral band resonances.
@@
Another noteworthy result in Ref. [397] is the demonstration of the generalized Brillouin zone using the Laplace transform rather than the Fourier transform applied to the measured voltage distribution. The non-Hermiticity renders the wavevector complex, making it infeasible to obtain the Brillouin zone directly from the Fourier transform. However, Wu et al. utilized the Laplace transform method to effectively recover the generalized Brillouin zone, where different onsite admittances correspond to different contours of the band structure. Such site-resolved field distributions are highly accessible in topolectrical circuits, which is crucial for studying higher-dimensional lattices. For example, in another work [401], Wu et al. observed degeneracy conversions, including Weyl point-to-nodal line conversion and nodal line-to-entangled nodal line conversion.
@@ IV.3 Hyperbolic lattices and their realizations in electrical circuits
Lately, there has been much interest in the band structure and topological properties of hyperbolic lattices, which can be regarded as a vast generalization of regular lattices in negatively curved space [405, 406, 407, 403, 408, 409, 410, 411, 412, 413, 414, 415]. In 2D Euclidean (flat) space, regular -gons can be tessellated such that every vertex contains branches, if . This leaves only three possibilities corresponding to the triangular, square and hexagonal lattices respectively. Relaxing this constraint to results in an infinite variety of so-called hyperbolic lattices, which however need to be embedded in negatively curved space 111The opposite constraint gives the five possibilities , respectively corresponding to the five platonic solids, i.e., icosahedron, dodecahedron, octahedron, cube and tetrahedron, all tiling the sphere.. While this requirement of negatively curved space embedding makes them challenging to realize in most metamaterial platforms, electrical circuit connections can describe any desired graph structure, and thus emerge as promising physical platforms for hyperbolic lattices. This negative curvature fundamentally alters the propagation of modes, enriching the behavior of fundamental phenomena such as topological propagation [417]. Due to the relative proliferation of boundary terminations, a hyperbolic lattice structure also alters the nature of the non-Hermitian bulk-boundary correspondences in the presence of asymmetric lattice hoppings [418]. The ability of electrical circuits to simulate lattices with high coordination numbers is also helpful in the realization of higher dimensional lattices [419, 420, 421, 114, 50, 129, 105, 130].
@@
where is the inductance and the phase is controlled by adjusting the external voltages of the analog multipliers. The diagonal entries introduce a constant offset in the admittance spectrum, while the off-diagonal terms generate the controllable complex phase between ports and .
A key feature of finite hyperbolic lattices is that the number boundary sites scales exponentially with the radius (i.e., number of lattice generations), reminiscent of fractal lattices [422]. As such, unlike in Euclidean lattices, boundary sites can dominate the Hilbert space and greatly alter bulk-boundary correspondences, even in Hermitian setups. Zhang et al. [403] demonstrated the exponential proliferation of zero modes in a hyperbolic circuit [Fig. 8c]. These topological zero modes give rise to a large number of impedance resonances at the edge nodes, while no resonances occur in the bulk nodes.
This boundary sensitivity is reflective of the small-world connectivity of a hyperbolic lattice, a property that has inspired their application in holographic duality [423, 424, 425, 426, 427, 428, 429]. The high connectivity also gives rise to heightened susceptibility to lattice defects or perturbations. Pei et al. [404] experimentally constructed a hyperbolic Haldane circuit and showed that a single defect in the bulk can induce boundary-dominated one-way propagation. By introducing polygonal bulk defects into the Haldane hyperbolic lattice, robust boundary propagation was observed, even when induced by a single point defect in the bulk [Fig. 8d]. While the presence of zero modes was verified through impedance responses, the one-way propagation of topological edge states was demonstrated using dynamical voltage analysis.
@@ IV.4 Simulations of other topological and quantum systems
Electrical circuits are also well suited for simulating quantum media, both of conventional topological materials and more unconventional setups such as [401, 430, 431, 432]. Described below are some experimental demonstrations, particularly those of models that are hard to simulate on other platforms.
One class of models that are particularly amenable to electrical circuit realizations are the square-root models. Given a parent Hamiltonian, the square-root of it is a lattice model in which two successive hoppings yield the same connectivity structure as the parent model (up to an unimportant constant shift) [445, 446, 447, 448, 449, 450, 451, 452, 453, 454, 455, 456, 457, 458]. Mathematically, they may look contrived and complicated, but they inherit various properties from their parent Hamiltonians, such as midgap topological modes [459]. The number of modes, including topological modes, is doubled in these lattices [460]. Electrical circuits are suitable physical platforms due to their versatile connectivity, and Song et al. [433] constructed a 3D square-root higher-order Weyl semi-metal by stacking 2D square-root lattices along the third dimension [Fig. 9a]. From the admittance spectrum, non-zero mid-gap topological hinge states were observed.
@@
Beyond band structures, 2D or 3D electrical circuit arrays can also simulate few-body physics and other phenomena. By mapping the 2-photon degrees of freedom onto a 2D lattice [48, 317, 467], Zhou et al. [434] simulated interaction-induced flat-band localizations and topological edge states, tuning the effective interaction strength through circuit grounding. By measuring impedance responses and voltage dynamics in the time domain, they were able to resolve two-boson flat bands and topological edge states [Fig. 9c]. On a different note, Yang et al. [435] simulated Wilson fermions—a lattice discretization conventionally used in the context of lattice QCD to avoid fermion doubling—in a circuit lattice, tuning circuit parameters to achieve different Wilson fermion states. Through impedance measurements, they reconstructed the chiral edge domain wall separating two circuits with contrasting fractional Chern numbers, and observed localized corner states arising from two Chern lattices with opposite chiralities [Fig. 9d]. Wang et al. [436] realized the elusive Hopf insulator in a 3D electrical circuit systematically implemented with so-called pseudospin modules. Although this lattice required long-range spin-orbit couplings, the pseudospin modules aided the realization of Hopf bands, including both surface and bulk bands [Fig. 9e].
Electrical circuits have also been used to simulate non-Abelian systems, in which physical outcomes depend on the sequence of operations, i.e., twists [468, 469, 470, 471, 472, 473]. Various recent experiments have attempted to classically simulate operations that are mathematically equivalent to that experienced by anyonic quasiparticles [474, 475, 476, 477, 478, 479]. Wang et al. [437] demonstrated non-Abelian Anderson localization using a 2D electrical circuit that represents the quasiperiodic AAH model with non-Abelian gauge fields. Both localization and delocalization behaviors were observed through site-resolved impedance spectra and voltage dynamics [Fig. 9f] (see also the experimental realization of non-Abelian inverse Anderson transition in Ref. [480]).
Circuits are also particularly suitable for demonstrating the effects of lattice defects, since nodes are connected at will. Xie et al. [438] experimentally demonstrated disclination states and their correspondence to real-space topology. Localized topological states at the disclinations—defects in rotational symmetry [481]—were observed in impedance maps and voltage distributions at the lattice nodes [Fig. 9g] (see also Refs. [482, 483, 484]).
@@
Since electrical circuits commonly involve resistive and active elements, they can readily demonstrate how non-Hermiticity interplays with topology without much additional complication. Xie et al. [440] demonstrated the non-Hermitian version of the graphene lattice, which exhibits two non-Hermitian Dirac cones, one experiencing amplification and the other experiencing decay. Site-resolved band measurements reveal unidirectional kink states [493, 494] with dissimilar profiles at the two valleys [Fig. 9j]. Involving asymmetrical real next-nearest-neighbor couplings effectively results in a non-Hermitian single Dirac cone, leading to valley polarization [495, 496, 497, 498] with a large bandwidth [see also Ref. [499]].
The introduction of non-Hermiticity can also greatly alter the locality of topological modes. While topological modes are typically strongly confined to the boundaries [30], non-Hermitian skin pumping in the opposite direction can delocalize them while preserving their eigenspectrum, in this case mid-gap zero-eigenvalues [500]. These delocalized topological states, still protected by lattice symmetries, can extend through the bulk, enabling their robustness to be harnessed not only at the boundaries but also within the bulk [501, 502]. Lin et al. [441] demonstrated these extended states in 1D, 2D, and 3D circuits by tuning the non-reciprocal coupling strength [Fig. 9k]. At a critical strength determined by both the topological and skin localization length, the topological modes become extended, a phenomenon that was also studied in Refs. [503, 504, 38, 505, 506].
@@
Topolectrical circuits can simulate not only classical systems, but also quantum setups [511, 512, 513]. By mapping the Hilbert space of few-qubit systems to the nodes of suitably designed circuits, Zou et al. [443] demonstrated [Fig. 9m] the mathematical equivalences of Majorana-like edge states, T-junctions, two-qubit unitary operations, and Grover’s search algorithm. Through their resistor-capacitor-based circuit with segmented fixed resistances, braiding processes can be simulated without significant loss. Zhang et al. [444] demonstrated another example of quantum algorithm simulation using classical circuits. They simulated a two-player quantum game based on AND-OR tree structures [Fig. 9n]. Relays were employed to disconnect each node simultaneously after applying the initial voltage, allowing the system to evolve from the initial state.